OpenStudy (anonymous):

solve 15t^2 -15t - 30 =0

OpenStudy (bahrom7893):

t^2 - t -2 = 0

OpenStudy (bahrom7893):

ill let nenad finish haha

OpenStudy (nenadmatematika):

15(t^2-t-2)=15t(t+1)(t-2)=0 so t=-1 or t=2 and thank you mr bahrom :D